### v1

# Language Model Evolutionary Algorithms for Recommender Systems: Benchmarks and Algorithm Comparisons

Abstract
In the evolutionary computing community, the remarkable language-handling capabilities and reasoning power of large language models (LLMs) have significantly enhanced the functionality of evolutionary algorithms (EAs), enabling them to tackle optimization problems involving structured language or program code. Although this field is still in its early stages, its impressive potential has led to the development of various LLM-based EAs. To effectively evaluate the performance and practical applicability of these LLM-based EAs, benchmarks with real-world relevance are essential. In this paper, we focus on LLM-based recommender systems (RSs) and introduce a benchmark problem set, named RSBench, specifically designed to assess the performance of LLM-based EAs in recommendation prompt optimization. RSBench emphasizes session-based recommendations, aiming to discover a set of Pareto optimal prompts that guide the recommendation process, providing accurate, diverse, and fair recommendations. We develop three LLM-based EAs based on established EA frameworks and experimentally evaluate their performance using RSBench. Our study offers valuable insights into the application of EAs in LLM-based RSs. Additionally, we explore key components that may influence the overall performance of the RS, providing meaningful guidance for future research on the development of LLM-based EAs in RSs.
Index Terms:
Evolutionary algorithm, large language models, recommender systems, multiobjective optimization.
I Introduction
Evolutionary algorithms (EAs), inspired by natural selection and genetic principles, constitute a powerful class of optimization techniques [1]. Unlike traditional methods, EAs do not rely on gradient information, making them highly suitable for tackling complex black-box optimization problems that commonly arise in scientific [2] and engineering [3] fields. Moreover, EAs are particularly effective for multiobjective optimization problems, where multiple conflicting objectives require simultaneous optimization [4]. Their population-based search strategy, which leverages implicit parallelism, allows EAs to explore diverse solution spaces within a single optimization run. This approach yields a set of optimized solutions—referred to as nondominated or Pareto optimal solutions—that align with the Pareto front (PF), representing the best achievable performance trade-offs in the objective space. Meanwhile, diversity-preserving mechanisms within EAs further enable the discovery of solutions that are well-distributed along the PF, allowing decision-makers to visualize the trade-offs and select the most preferred solution(s) a posteriori [5].
Large language models (LLMs)[6], such as ChatGPT[7], have significantly advanced language processing technology, exhibiting exceptional performance in natural language processing tasks and fostering transformative applications across fields such as mechanical engineering [8], management [9], and chemistry [10]. This advancement has created promising opportunities to enhance EAs [11]. On one hand, the black-box optimization capabilities of EAs are well-suited for LLM-based applications. Due to the high computational demands of LLM training and inference, deployment is largely restricted to corporations like OpenAI, Google, and Baidu, which often limit access to model parameters for security and privacy, providing only API-based querying [11]. Consequently, users often treat LLMs as black-box systems, and EAs, as classic and effective black-box optimization algorithms, present a promising approach for refining LLM applications [12]. On the other hand, the generative capabilities of LLMs can help expand the functional scope of traditional EAs, which typically represent solutions as low-dimensional continuous or discrete vectors [13]. Integrating LLMs enables EAs to process language and code, allowing them to tackle complex tasks such as prompt optimization [14], neural architecture search [15], and code generation [16]. Consequently, recent research has increasingly explored hybrid approaches that combine EAs with LLMs, applying these techniques to both traditional optimization challenges and novel application domains.
Recommender systems (RSs) play an increasingly crucial role on modern online platforms [17], and the emergence of LLMs also has introduced new opportunities for their development and enhancement [18]. In the early stages, language models such as BERT [19] were used to embed language elements from the recommendation process into suitable vectors, supporting the training of more accurate recommendation models [20]. As LLMs evolved, pre-training [21] and fine-tuning [22] paradigms were developed, leading to the creation of even more powerful recommendation performance [23]. Recently, with the advent of models like ChatGPT [7], the advanced language processing capabilities of LLMs enable the creation of robust RSs through simple prompts [24, 25, 26] or in-context learning [27, 28].
In prompt-based LLM-driven RSs, the design of prompts plays a pivotal role in shaping system performance [26]. As a result, finding the optimal prompt becomes a critical priority when developing such systems. Additionally, in RSs, it is essential for recommendations not only to align with user-specific interests but also to consider factors such as diversity and fairness to enhance overall user satisfaction [29, 30]. Multiobjective optimization serves as an ideal theoretical framework for this purpose [31]. By leveraging this approach, the system can provide users with a set of Pareto-optimal solutions that balance criteria like recommendation accuracy, diversity, and fairness, thus allowing users to select options that best meet their personalized preferences. Whether addressing black-box prompt optimization or multiobjective optimization, EAs have demonstrated remarkable capabilities. Consequently, EAs hold significant potential for application in LLM-based RSs. However, to the best of our knowledge, there has no research, both in terms of benchmark problems and algorithms, on applying EAs to LLM-based RSs. To bridge this gap, this paper explores the application of language model based EAs in RSs. The key contributions of this paper are mainly focus on the follow aspects:
  • We introduce a benchmark problem set, named RSBench, specifically designed for EAs on LLM-based RSs. RSBench focuses primarily on session-based recommendation [26], which aims to predict next promising items for interaction based on short, anonymous behavior sessions. The goal is to discover a set of Pareto optimal prompts that can guide the recommendation process, allowing users to freely choose recommendations based on their preferences regarding accuracy, diversity, and fairness, balancing two or three of these aspects.
  • A set of LLM-based multiobjective EAs are developed to solve the RSBench problems. Following the frameworks of three representative multiobjective EAs—NSGA-II, MOEA/D, and IBEA—LLM-based operations are integrated into these classical algorithms to support their application in solving RSBench problems. Their performance is then evaluated experimentally, providing valuable insights and guidance for the application of EAs in LLM-based RSs.
  • We further investigate key components that may impact the performance of the final RS, including the effectiveness of the LLM-based initialization and the LLM-based offspring generation with the supporting of the crossover and mutation operators. These investigations are expected to provide valuable insights and guidance for future research on LLM-based EAs in RSs.
  • Finally, with regard to LLM-based EAs for RSs, we discuss promising research topics and pressing issues that warrant future exploration, providing guidance for subsequent research.
The remainder of this paper is organized as follows: Section II provides a review of related work and background concepts. Section III introduces RSBench. In Section IV, we present the three LLM-based multiobjective EAs, based on the classical NSGA-II, MOEA/D, and IBEA frameworks. Section V discusses the systematic experimental investigations. In Section VI, we explore promising research directions based on RSBench and highlight pressing issues for future exploration. Finally, Section VII concludes the paper.
II Preliminaries
II-A Evolutionary Multiobjective Optimization
Without loss of generality, a multiobjective optimization problem can be stated as:
$\max:$ $\textbf{F}(\textbf{p})=\{F_{1}(\textbf{p}),\ldots,F_{m}(\textbf{p})\}$ (1)
s.t. $\textbf{p}\in\mathcal{X}$
where $F_{i}(\textbf{p}),(i\in\{1,\ldots,m\})$ is the $i$th objective function, $m$ is the number of objectives, p is the decision variable which is considered as a prompt in this paper, $\mathcal{X}$ is the decision space (or search space). Key concepts associated with the formulation in (1) are as follows [4]:
  • Pareto Dominance: For decision vectors $\textbf{p}_{a}$ and $\textbf{p}_{b}$, if $\forall i\in\{1,2,\ldots,m\}$, $F_{i}(\textbf{p}_{a})\geq F_{i}(\textbf{p}_{b})$ and $\exists j\in\{1,2,\ldots,m\}$, $F_{j}(\textbf{p}_{a})>F_{j}(\textbf{p}_{b})$, $\textbf{p}_{a}$ is said to Pareto dominate $\textbf{p}_{b}$.
  • Pareto Optimal Solution: If no decision vector in $\mathcal{X}$ Pareto dominates $\textbf{p}_{a}$, then $\textbf{p}_{a}$ is a Pareto optimal solution.
  • Pareto Set: The set of all Pareto optimal solutions forms the Pareto set in decision space.
  • Pareto Front: The image of the Pareto set in the objective space forms the Pareto front.
EAs, with their inherent parallelism in population-based search, have long been instrumental in solving multiobjective optimization problems, as noted in [32]. Multiobjective EAs are typically classified into three main categories: domination-based methods [33, 34], decomposition-based methods [35, 36], and indicator-based methods [37, 38]. Over recent decades, these three types of multiobjective EAs have consistently demonstrated strong performance across various multiobjective optimization tasks and have been extended to address many-objective optimization problems [39], constrained multiobjective optimization problems [40, 41], and multiobjective combinatorial optimization problems [42]. Distinct from previous studies, this paper further explores the integration of LLMs within these classical EA frameworks, examining their potential applications in RSs.
II-B Large Language Models in Evolutionary Computation
The integration of LLMs and evolutionary computation can be categorized into two primary directions [11]. The first direction focuses on utilizing LLMs to enhance EAs, thereby directly addressing optimization challenges or achieving improved performance on various mathematical programming problems [43]. These methods are founded on the premise that the extensive knowledge acquired by LLMs can guide EAs in developing more effective and efficient optimization frameworks. One of the earliest works in this domain is OPRO [44], which adapts LLMs to solve optimization problems by generating promising solutions for combinatorial optimization challenges, such as the traveling salesman problem. Liu et al. [43] propose LMEA, which leverages LLMs for crossover and mutation operations and constructs prompts in each generation to assist the LLM in selecting parent solutions from the current population. In addition to these methods, there is ongoing research investigating the effectiveness of various LLM-based components, including crossover/mutation [13], surrogate models [45], and their extension to constrained optimization [46], multiobjective optimization [47] and/or multitask optimization [48].
The second direction involves leveraging the powerful black-box optimization capabilities of EAs to address optimization problems within LLM-based application scenarios [49, 50, 51]. One prominent application area for EAs is prompt engineering [52], where various studies have explored the use of EAs to identify optimal prompts, represented as both textual instructions and embedded vectors, to guide LLMs in executing complex language tasks [53, 12, 54]. Another significant application is code generation, which benefits from the robust code generation capabilities of LLMs alongside the optimization strengths of EAs to produce high-quality code [55]. Recently, several EA-LLM-based methods for code generation have been proposed [50, 56]. Additionally, with the support of LLMs’ reasoning and coding abilities, EAs have also been effectively utilized for neural network architecture search, yielding impressive results [51, 57].
Despite the increasing number of studies combining EAs and LLMs to achieve various objectives, to the best of our knowledge, no research has yet applied these methodologies to RSs. In contrast to prior works, this paper primarily focuses on the application of EAs in the context of LLM-based RSs.
II-C Recommender Systems in Evolutionary Computation
EAs are well-suited for RSs due to their ability to handle multiobjective optimization, aligning with the multi-metric nature of recommendation tasks. Geng et al.[58] framed recommendation as a multiobjective optimization problem, proposing NNIA-RS, which utilizes a nondominated neighbor immune algorithm as an optimizer. Similarly, Cui et al.[59] applied multiobjective EAs to balance precision and diversity in recommendations. Beyond these approaches, recent advances, including many-objective optimization [60], evolutionary clustering [61], and hybrid methods combining EAs with neural networks [62], have further advanced EA applications in RSs. However, these methods are limited by their reliance on small-scale user-item interactions and struggle to integrate session-based natural language features, constraining their scalability and applicability in real-world contexts [63].
Recently, LLM-based RSs have demonstrated strong potential for generating context-aware, personalized recommendations, often requiring multiobjective optimization to incorporate human-centered, customized preferences [18]. Nevertheless, limited studies have explored the application of multiobjective EAs within LLM-based recommendation frameworks.
III RSBench: The Proposed Benchmark
III-A Description of the Recommendation Task
Figure 1: Examples of user sessions with various intents [26].
Figure 2: The example of the LLM-based RS.
In this paper, we focus on session-based recommendation [26], which aims to predict the next item a user will interact with based on short, anonymous behavior sessions. This approach is grounded in the observation that different sessions may reveal distinct user intents. In Fig. 1, we illustrate two real sessions from the Amazon Electronics dataset [64]. The first session, shown in (a), reflects a single primary intent—shopping for a laptop and related accessories—whereas the second session, shown in (b), demonstrates two major intents: shopping for laptop protectors and camera accessories. A key challenge in session-based recommendation is accurately inferring the user’s intent based on their interactions to provide relevant product suggestions. Leveraging the powerful reasoning capabilities of LLMs, the RS has the potential to more effectively capture latent user intents within a session, thereby improving recommendation performance [26]. This paper focuses on the development and evaluation of an LLM-assisted session-based RS.
The RS considered in this paper consists of three key components:
  • Recommendation Prompt: The recommendation prompt describes the current task, guiding the LLM to generate appropriate recommendations for the customer.
  • Session Interactions: For each customer, a set of session interactions is provided to the LLM. These interactions allow the LLM to infer the customer’s intent, based on their behavior within the session.
  • Candidate Set: A candidate set of products is supplied to the RS. The system then ranks these products, recommending the top-ranked items to the customer.
For clarity, we provide an example on the recommendation of the movie in Fig. 2 to illustrate how the proposed RS operates. In this example, the recommendation prompt shown is provided to the LLM, along with the session interactions and the candidate set. The LLM processes this information to generate a ranked list of items from the candidate set. The top-ranked items, specifically the top-5, are then recommended to the customer.
III-B RSBench Optimization Problems
In this subsection, we introduce the details of RSBench. First, we define the decision variables and objective functions used in RSBench problems. Then, we present the nine multiobjective instances, which serve as a recommended set of benchmark problems for investigating the performance of LLM-based EAs in RSs.
III-B1 Decision Variables
As discussed in Section III-A, the LLM’s recommendation is primarily guided by the recommendation prompt. Consequently, the quality and structure of the prompt have a significant impact on the performance of the RS. To maximize the system’s performance, it is crucial to identify optimal recommendation prompts. Therefore, in the RSBench problem set, the recommendation prompt, denoted as p, is treated as the decision variable.
III-B2 Objective Functions
Several indicators can be employed to evaluate the performance of the RS from various perspectives. In this paper, we primarily utilize the following three indicators to assess the RS, using them as objective functions.
  • Recommendation Accuracy: Recommendation accuracy is used to assess how effectively the RS identifies the customer’s intentions and suggests the desired items [65]. An empirical estimate of recommendation accuracy can be calculated using the existing ground truth dataset. Given a dataset $\mathcal{D}=\{(\textbf{X}_{s}^{(l)},\mathcal{Y}_{c}^{(l)},\textbf{y}_{t}^{(l)})\}_{l=1}^{N}$, where $\textbf{X}_{s}^{(l)}$ is a sequence of session, $\mathcal{Y}_{c}^{(l)}$ is a set of candidate items, and $\textbf{y}_{t}^{(l)}\in\mathcal{Y}_{c}^{(l)}$ is a ground truth target item selected by the customer, we employ the following function to estimate recommendation accuracy
    $F_{acc}(\textbf{p})=\frac{1}{N}\cdot\sum_{l=1}^{N}\mathbb{I}\left[R_{llm}(\textbf{y}_{t}^{(l)}|\textbf{X}_{s}^{(l)},\mathcal{Y}_{c}^{(l)},\textbf{p})\leq K\right].$ (2)
    In (2), $\mathbb{I}[\cdot]$ is an indicator function, and $R_{llm}(\textbf{y}_{t}^{(l)}|\textbf{X}_{s}^{(l)},\mathcal{Y}_{c}^{(l)},\textbf{p})$ is the predicted ranking of $\textbf{y}_{t}^{(l)}$ provided by the LLM with the input of $\textbf{X}_{s}^{(l)}$, $\mathcal{Y}_{c}^{(l)}$, and the recommendation prompt p. From (2), we observe that recommendation accuracy considers a recommendation correct when the target item is among the top-K-ranked items in the candidate set. The accuracy rate is then calculated as the indicator to evaluate the performance of the recommendation.
  • Recommendation Diversity: In many cases, the suggested items for customers may contain similar or overlapping information [29]. However, in real-world RSs, it is equally important to offer a diverse set of recommendations to keep users engaged by introducing novel content they may not have explicitly searched for but might still enjoy [66]. To achieve this, we aim to provide recommendations that exhibit a certain degree of diversity [67]. In this context, the information of an item is represented by its category, and the diversity is calculated as follows
    $F_{div}(\textbf{p})=\frac{1}{N}\cdot\sum_{l=1}^{N}\frac{|\cup_{k=1}^{K}C[T^{k}_{llm}(\textbf{X}_{s}^{(l)},\mathcal{Y}_{c}^{(l)},\textbf{p})]|}{\sum_{k=1}^{K}\left|C[T^{k}_{llm}(\textbf{X}_{s}^{(l)},\mathcal{Y}_{c}^{(l)},\textbf{p})]\right|}$ (3)
    where $T^{k}_{llm}(\textbf{X}_{s}^{(l)},\mathcal{Y}_{c}^{(l)})$ denotes the item ranked $k$ by the LLM-based RS, $C[\cdot]$ represents the category set associated with an item, and ||$\cdot$|| ⋅ | denotes the size of a set. The formula first calculates the ratio of the number of unique categories among the top-$K$ items to the sum of the categories of each top-$K$ item for each recommendation sample. It then computes the average ratio across all N recommendation samples. This formula reflects diversity by assigning smaller ratios to cases with high category overlap among top-$K$ items, which implies reduced diversity. Hence, higher values of $F_{div}$ indicate greater diversity in the recommendations.
  • Recommendation Fairness: Recommendation fairness is another important indicator considered in this paper [30]. In RSs, biases can easily arise. For instance, the system may favor recommending popular items that have been available for a longer period, as these products typically exhibit higher click-through and purchase rates. Newly launched items, however, may lack sufficient historical data, causing them to be overlooked, even if they better meet customer needs [68]. Therefore, ensuring fairness in recommendations is crucial. In this paper, we calculate recommendation fairness based on the click-through rates of the items [68]. First, we divide all items into two sets: $\mathcal{H}$, which contains the top 20% of items based on click-through rates, and $\mathcal{L}$, which includes the remaining 80%. The recommendation fairness is then calculated as follows:
    $F_{fair}(\textbf{p})=\frac{1}{N}\cdot\sum_{l=1}^{N}\frac{\sum_{k=1}^{K}\mathbb{I}[T_{llm}^{k}(\textbf{X}_{s}^{(l)},\mathcal{Y}_{c}^{(l)},\textbf{p})\in\mathcal{L}]}{K}.$ (4)
    The formula calculates the ratio of items with lower click-through rates within the top-$K$ items for each recommendation sample, then computes the average ratio across $N$ samples. A higher $F_{fair}$ value indicates a greater tendency to recommend items with lower click-through rates, thus reflecting improved fairness in the recommendation outcomes.
Remark: It should be noted that all the objectives introduced above are maximization objectives.
III-B3 Instances of RSBench
TABLE I: The Statistical Characteristics of Datasets. “Density Indicator” Refers to the Average Frequency of Each Item Appearing in the Dataset.
Items Sessions Ave. Session Length Density Indicator
ML-1M 3416 784860 6.85 1573.86
Games 17389 100018 4.18 24.04
Bundle 14240 2376 6.73 1.12
Based on the definitions of the decision variables and objective functions provided above, we introduce nine instances of the RSBench problem set. To obtain the objective function values, datasets containing customer consumption and click data are required. In this benchmark, we use three real-world datasets from different domains. Specifically, MovieLens-1M (ML-1M) [^1] https://grouplens.org/datasets/movielens/ is a movie dataset containing customer ratings for various films. The “Games” dataset is a subcategory from the Amazon dataset [64], which includes customer ratings for video games. The “Bundle” dataset [69] consists of session data from three Amazon subcategories: Electronics, Clothing, and Food. The statistical features of these datasets are summarized in Table I. For the ML-1M and Games datasets, we chronologically order the rated items for each user and divide them into sessions by day. All datasets are split into training and validation sets with a ratio of 8:2.
TABLE II: The Settings of the RSBench Problems. “Dataset” means the dataset used in the benchmark problem, “Batch Size” represents the number of selected samples used for estimating the objective function, $m$ denotes the number of objectives, and “Objective Functions” are the evaluation indicators considered in the benchmark problem.
Dataset Batch Size $m$ Objective Functions
RSBench-1 ML-1M 10 2 $F_{acc},F_{div}$
RSBench-2 Games 10 2 $F_{acc},F_{div}$
RSBench-3 Bundle 10 2 $F_{acc},F_{div}$
RSBench-4 ML-1M 10 2 $F_{acc},F_{fair}$
RSBench-5 Games 10 2 $F_{acc},F_{fair}$
RSBench-6 Bundle 10 2 $F_{acc},F_{fair}$
RSBench-7 ML-1M 10 3 $F_{acc},F_{div},F_{fair}$
RSBench-8 Games 10 3 $F_{acc},F_{div},F_{fair}$
RSBench-9 Bundle 10 3 $F_{acc},F_{div},F_{fair}$
Using these three datasets, the settings for the nine RSBench instances are presented in Table II. All RSBench problems are multiobjective optimization problems. Note that the LLM-based RS requires continuous online interaction with the LLM during the recommendation process. As this interaction process is time-consuming, it is impractical to evaluate the objectives using the entire training set. Therefore, when calculating the objective function values, we randomly select a batch of samples from the training set and compute the results based on this batch. As shown in Table II, the batch size for all RSBench problems in this paper is set to 10. However, users of this benchmark may choose a larger batch size for more accurate objective function estimation, although this may result in longer evaluation times. Developing a more efficient and time-saving evaluation strategy is a promising direction for further exploration, though it is beyond the scope of this paper.
IV Language Model Multiobjective Evolutionary Algorithms for RSBench
In this section, we present the details of the LLM-based EAs employed for RSBench in this paper. First, we describe two fundamental operations utilized in the algorithms: the initialization of the recommendation prompt population and the generation of offspring recommendation prompts. Next, we provide a detailed explanation of the three LLM-based EAs.
Figure 3: The example recommendation prompt designed following the chain-of-thought.
Figure 4: The prompt for the initialization of the population.
IV-A Basic Operations
IV-A1 Initialization of the Recommendation Prompt Population
While the robust capabilities of LLMs enable them to excel at various downstream tasks, their performance is highly dependent on the quality of the prompts provided. Over the past year, numerous prompt structures, such as chain-of-thought [70] and tree-of-thought [71], have been proposed to improve LLM performance. In this paper, to expedite the discovery of high-quality prompts, we introduce a well-designed recommendation prompt as the example based on the chain-of-thought approach. The details of this example recommendation prompt are presented in Fig. 3. The initial population of recommendation prompts is subsequently generated using the LLM based on the example recommoned prompt, with the initialization prompt shown in Fig. 4. In this manner, $N$ recommendation prompts are produced to form the initial population.
IV-A2 Offspring Recommendation Prompts Generation
Figure 5: The prompt for the offspring generation.
In traditional EAs, crossover and mutation operators are used to generate offspring solutions. However, these operators are typically designed for continuous or discrete vectors and are not well-suited for handling natural language, such as recommendation prompts. Recent studies have demonstrated that, with appropriate prompts, LLMs can themselves serve as effective crossover and mutation operators for natural language processing tasks [12]. Following this approach, we employ an LLM-based crossover and mutation mechanism in this work. The prompt used for the LLM-based crossover and mutation operator is illustrated in Fig. 3. By providing two parent recommendation prompts to the LLM-based operator, an offspring recommendation prompt is generated. This LLM-based operator thus facilitates the generation of an offspring population from the parent population.
IV-B Main Frameworks
By incorporating the operations introduced in Section IV-A into multiobjective evolutionary frameworks, LLM-driven multiobjective EAs can be developed. In this paper, we consider three representative EA frameworks: NSGA-II, MOEA/D, and IBEA. By integrating LLM-based operations, we derive three LLM-based EAs, denoted as LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA, respectively.
IV-B1 LLM-NSGA-II
Algorithm 1 LLM-NSGA-II
[…]
0:  The nondominated recommendation prompt set $\mathcal{P}$;
1:  Generate the initial population $\mathcal{P}=\{\textbf{p}_{1},\ldots,\textbf{p}_{N}\}$ by using the initial operations introduced in Section IV-A1;
2:  Evaluate the initial population by using the objective functions;
3:  $t\leftarrow 0$
4:  while $t<t_{max}$ do
5:     Generate an offspring population of recommendation prompts $\mathcal{Q}=\{\textbf{q}_{1},\ldots,\textbf{q}_{N}\}$ according to the operations introduced in Section IV-A2;
6:     Evaluate the offspring population by using the objective functions;
7:     $\mathcal{P}\leftarrow\mathcal{P}\cup\mathcal{Q}$;
8:     Sorting the recommendation prompts in $\mathcal{P}$ according to the nondominated sorting and the crowd distance, and retain only the top $N$ recommendation prompts;
9:     $t\leftarrow t+1$;
10:  end while
11:  return  A population of nondominated recommendation prompts $\mathcal{P}$.
The LLM-NSGA-II follows the classical NSGA-II framework, with its pseudo-code presented in Algorithm 1. Initially, a population $\mathcal{P}$ of $N$ recommendation prompts is generated using the initialization operation described in SectionIV-A1, and their performance is evaluated using the defined objective functions. In each iteration, an offspring population $\mathcal{Q}$ is generated using the crossover operation introduced in Section IV-A2. The offspring prompts are then evaluated based on the same objective functions, and elitist selection is performed to choose $N$ recommendation prompts from the combined population $\mathcal{P}\cup\mathcal{Q}$ using nondominated sorting and crowding distance [33]. These selected prompts form the new population for the next iteration. Upon satisfying the termination conditions, the final population of nondominated recommendation prompts is returned.
IV-B2 LLM-MOEA/D
Algorithm 2 LLM-MOEA/D
[…]
0:  The nondominated recommendation prompt set $\mathcal{P}$;
1:  Generate the neighborhood set $\mathcal{B}_{i}$ for the weight vector $\textbf{w}_{i}$, where $i\in\{1,\ldots,N\}$;
2:  Generate the initial population $\mathcal{P}=\{\textbf{p}_{1},\ldots,\textbf{p}_{N}\}$ by using the initial operations introduced in Section IV-A1;
3:  Evaluate the initial population by using the objective functions;
4:  Initial the external population $\mathcal{E}$ and the reference point $\textbf{z}=(z_{1},...,z_{m})^{T}$;
5:  $t\leftarrow 0$
6:  while $t<t_{max}$ do
7:     for $i=\{1,\ldots,N\}$ do
8:        Randomly select two indexes $k$ and $l$ from $\mathcal{B}_{i}$, and generate a new prompt $\textbf{p}^{new}$ according to $\textbf{p}^{k}$ and $\textbf{p}^{l}$ by using the LLM-based offspring generation introduced in Section IV-A2;
9:        Evaluate $\textbf{p}^{new}$ by using the objective functions;
10:        Update the reference point z, the population $\mathcal{P}$ and the external population $\mathcal{E}$ following the MOEA/D framework;
11:     end for
12:     $t\leftarrow t+1$;
13:  end while
14:  return  An external population of nondominated recommendation prompts $\mathcal{E}$.
Decomposition-based evolutionary multiobjective optimization is a classical paradigm in multiobjective optimization. LLM-MOEA/D follows the decomposition-based method, with the pseudo-code presented in Algorithm 2. As in the MOEA/D framework, a neighborhood set $\mathcal{B}^{i}$ is initialized for each weight vector $\textbf{w}^{i}$. The initial population $\mathcal{P}$ is generated using the LLM-based initialization described in Section IV-A1, and all recommendation prompts in $\mathcal{P}$ are evaluated using the objective functions. Based on these evaluations, the external population $\mathcal{E}$ and the reference point z are initialized. During the iteration process, for each weight vector, an offspring recommendation prompt $\textbf{p}^{new}$ is produced using the LLM-based offspring generation method detailed in Section IV-A2. After evaluating $\textbf{p}^{new}$, the reference point z, the population $\mathcal{P}$, and the external population $\mathcal{E}$ are updated in accordance with the MOEA/D framework [36]. Once the termination conditions are satisfied, the external population $\mathcal{E}$ of nondominated recommendation prompts is returned.
IV-B3 LLM-IBEA
Algorithm 3 LLM-IBEA
[…]
0:  The nondominated recommendation prompt set $\mathcal{P}$;
1:  Generate the initial population $\mathcal{P}=\{\textbf{p}_{1},\ldots,\textbf{p}_{N}\}$ by using the initial operations introduced in Section IV-A1;
2:  Evaluate the initial population by using the objective functions;
3:  $t\leftarrow 0$
4:  while $t<t_{max}$ do
5:     Generate an offspring population of recommendation prompts $\mathcal{Q}=\{\textbf{q}_{1},\ldots,\textbf{q}_{N}\}$ according to the operations introduced in Section IV-A2;
6:     Evaluate the offspring population by using the objective functions;
7:     $\mathcal{P}\leftarrow\mathcal{P}\cup\mathcal{Q}$;
8:     Select $N$ recommendation prompts from $\mathcal{P}$ based on the $I_{(\epsilon)^{+}}$ indictor based environment selection, and forming a new generation of population;
9:     $t\leftarrow t+1$;
10:  end while
11:  return  A population of nondominated recommendation prompts $\mathcal{P}$.
The most classical indicator-based evolutionary algorithm, i.e., IBEA [38], is used to drive the evolution of LLM-IBEA. Similar to the previous two algorithms, it begins by initializing a population $\mathcal{P}$ using the LLM-based initialization process. The offspring population $\mathcal{Q}$ is then generated through the LLM-based crossover and mutation. After evaluating the offspring recommendation prompts, the $I_{\epsilon^{(+)}}$ indicator-based environmental selection [38] is applied to select $N$ recommendation prompts, forming a new population for the next iteration. Once the termination conditions are met, the final population of nondominated recommendation prompts is returned. The overall process of the LLM-IBEA is shown in Algorithm 3
TABLE III: HV results of LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA on RSBench Problems, averaged over 5 independent runs with diverse generated candidate sets.
Problems Training Set Validation set
LLM-NSGA-II LLM-MOEA/D LLM-IBEA LLM-NSGA-II LLM-MOEA/D LLM-IBEA
Average HV $\pm$ Std Average HV $\pm$ Std Average HV $\pm$ Std Average HV $\pm$ Std Average HV $\pm$ Std Average HV $\pm$ Std
RSBench-1 0.6606$\pm$0.0145 0.6652$\pm$0.0113 0.7088$\pm$0.0045 0.4978$\pm$0.0350 0.4877$\pm$0.0188 0.5186$\pm$0.0106
RSBench-2 0.5901$\pm$0.0333 0.6219$\pm$0.0270 0.6694$\pm$0.0386 0.4155$\pm$0.0198 0.3951$\pm$0.0254 0.4413$\pm$0.0570
RSBench-3 0.7637$\pm$0.0691 0.7578$\pm$0.0134 0.7448$\pm$0.0235 0.5308$\pm$0.0382 0.5122$\pm$0.0225 0.4302$\pm$0.0293
RSBench-4 0.4169$\pm$0.0284 0.5586$\pm$0.0411 0.5137$\pm$0.0413 0.3302$\pm$0.0592 0.3877$\pm$0.0217 0.3551$\pm$0.0338
RSBench-5 0.4143$\pm$0.0263 0.3615$\pm$0.0160 0.4171$\pm$0.0220 0.2765$\pm$0.0406 0.2433$\pm$0.0230 0.2400$\pm$0.0271
RSBench-6 0.2680$\pm$0.0273 0.2451$\pm$0.0211 0.2842$\pm$0.0231 0.1426$\pm$0.0271 0.1308$\pm$0.0165 0.1523$\pm$0.0244
RSBench-7 0.3511$\pm$0.0113 0.3914$\pm$0.0277 0.4116$\pm$0.0281 0.2254$\pm$0.0208 0.2311$\pm$0.0235 0.2142$\pm$0.0270
RSBench-8 0.2834$\pm$0.0238 0.2509$\pm$0.0079 0.2676$\pm$0.0250 0.1510$\pm$0.0239 0.1341$\pm$0.0151 0.1597$\pm$0.0229
RSBench-9 0.2580$\pm$0.0291 0.2724$\pm$0.0207 0.2652$\pm$0.0318 0.1186$\pm$0.0232 0.1225$\pm$0.0204 0.1250$\pm$0.0227
Figure 6: Pareto front approximation results provided by LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA on RSBench-1, RSBench-5, and RSBench-9 with the randomly sampled batch in training set. (a) LLM-NSGA-II results on RSBench-1. (b) LLM-MOEA/D results on RSBench-1. (c) LLM-IBEA results on RSBench-1. (d) LLM-NSGA-II results on RSBench-5. (e) LLM-MOEA/D results on RSBench-5. (f) LLM-IBEA results on RSBench-5. (g) LLM-NSGA-II results on RSBench-9. (h) LLM-MOEA/D results on RSBench-9. (i) LLM-IBEA results on RSBench-9.
V Experimental Studies
V-A Parameter Settings and Implementation Details
In our experiment studies, GPT-3.5 [7] is employed in both the RSBench and the LLM-based EAs. For the algorithms LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA, we set the maximum number of iterations to 20 and the population size to 10. In LLM-MOEA/D, 10 weight vectors are generated using the Riesz s-energy method [72], with a neighborhood size of 5. The environmental selection mechanisms for LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA adhere to those specified in the original papers [33, 36, 38]. In each iteration, a batch of 10 samples is randomly selected from the training set to evaluate objective function values. For each sample, 20 items are saved in the candidate set; if the target item ranks within the top 10 after re-ranking by the RS, the recommendation is deemed correct. The optimization process is repeated five times, each with a different random seed (i.e., 0, 10, 42, 625, and 2023) to generate diverse candidate sets. For the validation set evaluation, we randomly select 50 samples to estimate the objective functions, maintaining all other settings consistent with those used for the training set.
V-B Hypervolume Measure
We evaluated the algorithm quality using the hypervolume (HV) metric, which measures the size of the objective space region dominated by a set of solutions [73]. Since the objective values in the RSBench problems are maximization problems and confined to the region $[0,1]$, we use the point $(0,0)$ as the reference for two-objective problems (i.e., RSBench-1 to RSBench-6) and $(0,0,0)$ as the reference for three-objective problems (i.e., RSBench-7 to RSBench-9).
V-C Results of LLM-Based EAs on RSBench
Figure 7: Pareto front approximation results provided by LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA on RSBench-1, RSBench-5, and RSBench-9 with the randomly sampled batch in validation set. (a) LLM-NSGA-II results on RSBench-1. (b) LLM-MOEA/D results on RSBench-1. (c) LLM-IBEA results on RSBench-1. (d) LLM-NSGA-II results on RSBench-5. (e) LLM-MOEA/D results on RSBench-5. (f) LLM-IBEA results on RSBench-5. (g) LLM-NSGA-II results on RSBench-9. (h) LLM-MOEA/D results on RSBench-9. (i) LLM-IBEA results on RSBench-9.
TABLE IV: Consumed time of LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA on RSBench Problems, averaged over 5 independent runs with diverse generated candidate sets.
Problems LLM-NSGA-II LLM-MOEA/D LLM-IBEA
Average Consumed Time (seconds)
RSBench-1 11565.23 9583.45 10536.54
RSBench-2 14879.32 13439.38 14358.89
RSBench-3 28875.37 25673.86 21603.25
RSBench-4 12669.37 9498.68 10811.16
RSBench-5 13543.74 10783.86 13733.42
RSBench-6 24326.51 20826.94 24127.54
RSBench-7 11720.80 10425.90 14535.28
RSBench-8 17261.30 13521.75 14299.40
RSBench-9 16191.24 19020.70 20244.55
In this subsection, we assess the performance of LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA on RSBench problems. Table III presents the HV results achieved by these algorithms across multiple candidate sets generated with different random seeds. The results reveal that LLM-IBEA exhibits the best overall performance, achieving the highest HV values on five RSBench problems (RSBench-1, RSBench-2, and RSBench-5 to RSBench-7) with the training sets and on five problems (RSBench-1, RSBench-2, RSBench-5, RSBench-8, and RSBench-9) with the validation sets. LLM-NSGA-II and LLM-MOEA/D show comparable performance, with each achieving the highest HV results on two RSBench problems with both the training and validation sets. Specifically, LLM-NSGA-II performs best on RSBench-3 and RSBench-8 with the training set, and on RSBench-3 and RSBench-5 with the validation set. LLM-MOEA/D achieves top performance on RSBench-4 and RSBench-9 with the training set, and on RSBench-4 and RSBench-7 with the validation set.
Fig. 6 and 7 illustrate the objective-space approximate points for nondominated recommendation prompts generated by LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA on RSBench-1, RSBench-5, and RSBench-9 for both the training and validation sets. As shown, all three algorithms produce recommendation prompts that are well distributed across the objective space and display complementary dominance characteristics. This distribution enables users to choose high-quality recommendation prompts that best align with their personal preferences, thereby enhancing customization.
Additionally, Table IV presents the time each algorithm took to solve each problem. The results indicate that optimizing RSBench problems with LLM-based EAs is notably time-intensive, with each optimization run taking between approximately 2.5 and 8 hours. This high time consumption is primarily due to frequent interactions with the LLM for objective function evaluations in the RSBench problems. Both the inherent latency of LLM inference and the overhead from repeated communication contribute to the significant time costs associated with these evaluations.
V-D Effectiveness of the LLM-Based Initialization
Figure 8: Comparisons between LLM-IBEA and LLM-IBEA-wo-Init results. (a) HV results on RSBench-1 and RSBench-4 with the training sets. (b) HV results on RSBench-1 and RSBench-4 with the validation sets. (c) Pareto front approximations on RSBench-1 with the training set. (d) Pareto front approximations on RSBench-1 with the validation set. (e) Pareto front approximations on RSBench-4 with the training set. (f) Pareto front approximations on RSBench-4 with the validation set.
In Section IV-A1, we introduce an initialization strategy designed to generate the initial population of recommendation prompts for the three LLM-based EAs considered in this study. This strategy involves providing an example chain-of-thought recommendation prompt, which the LLM uses as a basis to generate the initial population. To assess whether this initialization strategy improves the performance of LLM-based EAs, we examine its impact using LLM-IBEA, the best-performing algorithm. Specifically, we create a variant, LLM-IBEA-wo-Init, which omits the example prompt during initialization, enabling a direct comparison of the two approaches. RSBench-1 and RSBench-4 are used as representative problems in this investigation.
Figure 8 presents the results of LLM-IBEA and its variant, LLM-IBEA-wo-Init, on the RSBench-1 and RSBench-4 problems. On the training set, as shown in Fig. 8, LLM-IBEA achieves higher HV results compared to LLM-IBEA-wo-Init for both problems. The approximating points in the objective space, depicted in Fig.8 and Fig. 8, reveal that LLM-IBEA demonstrates superior convergence. Additionally, the solutions provided by LLM-IBEA often include higher-accuracy recommendations, while LLM-IBEA-wo-Init lacks such high-accuracy solutions. On the validation set, similar trends are observed in Fig.8, Fig.8, and Fig. 8, where LLM-IBEA continues to outperform LLM-IBEA-wo-Init. This performance difference can be explained by the chain-of-thought example used in the initialization strategy, which exhibits strong recommendation accuracy and is closer to the PF of the optimization problem. By adapting the recommendation prompt based on this example, the initialization strategy generates an initial population that is nearer to the PF and includes recommendation prompts with high accuracy, thereby supporting faster convergence and improved solution quality.
V-E Effectiveness of the LLM-Based Offspring Generation
Figure 9: Comparisons between LLM-IBEA and LLM-IBEA-wo-C&M results. (a) HV results on RSBench-1 and RSBench-4 with the training sets. (b) HV results on RSBench-1 and RSBench-4 with the validation sets. (c) Pareto front approximations on RSBench-1 with the training set. (d) Pareto front approximations on RSBench-1 with the validation set. (e) Pareto front approximations on RSBench-4 with the training set. (f) Pareto front approximations on RSBench-4 with the validation set.
As described in Section IV-A2, the LLM-based EAs in this paper generate offspring recommendation prompts primarily through LLM-driven crossover and mutation operators. An important question to investigate is whether these LLM-driven crossover and mutation operators truly enhance algorithm performance on RSBench problems. To address this, we examine the performance of LLM-IBEA by comparing it with a variant named LLM-IBEA-wo-C&M, which omits the LLM-driven crossover and mutation operations. In LLM-IBEA-wo-C&M, offspring are generated directly based on all recommendation prompts stored in the population. The generation process is guided by the prompt: ‘Please generate a new prompt based on the following set: {population}. Provide a final prompt bracketed with <<<START>>> and <<<END>>>.’ We then evaluate the performance of both algorithms on RSBench-1 and RSBench-4.
Fig. 9 presents the results of LLM-IBEA and LLM-IBEA-wo-C&M on RSBench-1 and RSBench-4. For the training sets, as shown in Fig. 9, LLM-IBEA outperforms LLM-IBEA-wo-C&M on both problems in terms of HV results. Additionally, the approximated points in the objective space for LLM-IBEA tend to exhibit greater dispersion (i.e., Fig.9) and improved convergence (i.e., Fig.9), while the results from LLM-IBEA-wo-C&M show limited diversity and suboptimal convergence. On the validation set, similar phenomenon can be observed from Fig. 9, Fig. 9, and Fig. 9. The results from these experiments suggest that the LLM-driven crossover and mutation operations enhance solution diversity in the objective space. This improvement may stem from the LLM-driven crossover and mutation, which, unlike direct LLM-based conditional generation, allow the algorithm to explore a broader search space, leading to improved exploration capabilities and greater diversity among the resulting solutions.
VI What Can We Further Do Based on RSBench?
RSBench as a benchmark for RSs, providing strong evidence for the potential of EAs in LLM-based RSs. Looking ahead, RSBench is anticipated to catalyze further advancements in applying EAs within this field. In this section, we further present potential research points based on RSBench, with the goal of inspiring scholars in related disciplines to explore this promising area.
VI-A More Effective Objective Estimations
In RSBench, all objective function values are estimated using the Monte Carlo method, based on samples drawn from the training set. In our experimental studies, we randomly sample a batch of 10 samples from the training set to estimate each objective function value. However, it is clear that 10 samples may not be sufficient, and estimates based on this sample size may not accurately capture the quality of recommendation prompts. This limitation is reflected in our results, where the performance of optimized recommendation prompts on the test set is notably lower than on the training set. Increasing the sample size could mitigate this issue, but it would require more interactions with the LLM, significantly increasing the time and the token spent on evaluating each objective function. Consequently, designing more efficient, accurate, and time-saving objective function estimation methods is an area worth further exploration. Fortunately, RSBench can serve as a valuable platform for testing and evaluating these estimation techniques.
VI-B More Efficient Evolutionary Algorithms
In the search for optimal recommendation prompts, EAs rely on iterative, trial-and-error processes, requiring numerous objective function evaluations to identify high-quality prompts. In our experiments, we set the iteration count to 20 and the population size to 10, yielding a total of 200 evaluations. Each evaluation involves interaction with the LLM, resulting in considerable time costs. Reducing evaluation costs during the evolutionary process is therefore a key research challenge. LLM-based surrogate models [45] or evolutionary transfer/multitasking techniques [74, 75] offer promising directions. RSBench thus serves as a valuable benchmark for evaluating the efficiency of these approaches.
VI-C More Powerful Recommended Capabilities
In this paper, the considered RS has a simple structure, relying on a single recommendation prompt to generate recommendations based on user-provided information. However, there is considerable potential to enhance both the design of recommendation prompts and the overall system architecture. For instance, EAs could be employed to generate more sophisticated prompts, such as chain-of-thought [70] or even tree-of-thought [71] prompts, to improve system performance. Additionally, integrating techniques like in-context learning [76] or retrieval-augmented generation [77], with evolutionary computation optimizing relevant components, may make the LLM-based RS more robust. On such research topics, RSBench can serve as a flexible platform to support such expansions, enabling further exploration and innovation in this area.
VII Conclusion
To rigorously assess the performance and applicability of LLM-based EAs, real-world benchmark problems are crucial. This paper introduces RSBench, a benchmark problem set designed to evaluate LLM-based EAs in the context of recommendation prompt optimization, with a focus on session-based RS with the LLM. RSBench aims to identify a set of Pareto optimal prompts that guide the recommendation process, delivering recommendations that are accurate, diverse, and fair. Moreover, by integrating an LLM-based initialization strategy along with LLM-driven offspring generation strategies, we developed three LLM-based EAs—LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA—within established multiobjective evolutionary frameworks. Performance evaluations on RSBench reveal that LLM-IBEA achieves the best results across the benchmark problems. Furthermore, our analysis highlights the effectiveness of the LLM-based initialization and LLM-driven offspring generation in enhancing algorithmic performance. We also discuss future research directions that RSBench opens for LLM-based recommendation systems and EAs. We believe RSBench will support and drive future advancements in LLM-based EAs and RSs. Future work will focus on exploring more effective LLM-based EAs and recommendation methodologies, further advancing the application of evolutionary computing in the domain of recommendation systems.
Acknowledgment
The authors acknowledge the use of GPT-4o to identify improvements in the writing, prompts used with “Please refine and optimize the description of the following, and make it be suitable for a research paper”.